If 4 workers can paint 5 trucks 1 hour, how many will be needed to paint 200 trucks in 8 hours?
Thank You.

Answer by stanbon(75887)

If 4 workers can paint 5 trucks 1 hour, how many will be needed to paint 200 trucks in 8 hours?
If 4 workers can paint 5 trucks in 1 hour
4 workers = 5 trucks/ 1 hr.
Divide both sides by 4 to get
1 worker = (5/4) truck/ 1 hr.
200/(5/4)= 160
160 workers = 160(5/4) trucks/ 1hr
160 workers = 200 trucks/ 1 hr.
20 workers = 200 trucks/ 8hr.
Cheers,
Stan H.